OpenStudy (anonymous):

how do you differentiate 4^3^x^2?

OpenStudy (anonymous):

u bring the power behind the intiger then minus 1 ! for example if u have x^3 its gonna be 3x^2

OpenStudy (anonymous):

that might work for x^2, but im pretty sure you use the chain rule to solve this problem, because its a constant to a constant to a variable to a constant

OpenStudy (turingtest):

you are talikng about\[4^3\]to the \[x^4\]power, correct?

OpenStudy (turingtest):

sorry\[ x^2\] power

OpenStudy (anonymous):

Latex is failing to show that much to the power.

OpenStudy (anonymous):

yes 4^(3^(x^(2)))

OpenStudy (turingtest):

ok, notice that 4^3 is a constant 4^3=64 so we have \[64^{x^2}\]and\[{d \over dx}(a^x)=a^x \ln( a)\]with that and the chain rule you should be able to solve this

OpenStudy (anonymous):

Alright,it's not much difficult,just apply the chain rule repeatedly. Answer would be: \[ 2^{1+6^{x^2}} \times 3^{x^2} \times x \times \text{ln}(3) \text{ln}(4) \]

OpenStudy (anonymous):

when i use the chain rule should i make f(x)=x^2, then g(x)=64^x?

OpenStudy (jamesj):

The way colton has written the function means we can't 'contract' the 4^3. With y = 4^(3^(x^2)) then ln y = 3^(x^2) ln 4 and thus ln(ln y) = x^2 . ln 3 + ln(ln 4) Now differentiate both sides, using implicit differentiation on the left hand side.

OpenStudy (anonymous):

You may also check here: http://www.wolframalpha.com/input/?i=derivative+of+4^%283^%28x^%282%29%29%29

OpenStudy (turingtest):

if it's g(f(x)) then yes

OpenStudy (turingtest):

wow, I never anticipated such a difficult problem. I figure it had to be easier. that was clearly a mistake. ?

OpenStudy (anonymous):

im just in the beginning of calc 1, so i dont know implicit differentiation yet, but i believe that i can do 4^3 first. when i wrote the equation i was trying to make it clearer

OpenStudy (turingtest):

No James know what he's talking about and so does wolfram and colton, I was wrong, you cannot do that

OpenStudy (turingtest):

check the link colton posted

OpenStudy (anonymous):

foolformath's link has the equation written correctly

OpenStudy (jamesj):

If the function is f(x) = 4^(3^(x^2)) and you can't use implicit differentiation, the chain rule will do it. Let g(x) = x^2, h(x) = 3^x and j(x) = 4^x then f(x) = j(h(g(x))) and\[{df \ \over dx} = {dj \over dh}{dh \over dg}{dg \over dx}\]

OpenStudy (anonymous):

ok thanks

OpenStudy (jamesj):

This is a bit tricky. Try differentiating the function h(g(x)) first. And then layer on j